OpenStudy (anonymous):

please help! Prove the triangle exterior angle theorem The measure of each exterior angle of a triangle equals the sum of the measures of its two remote interior angles Given: <1 is an exterior angle of the triangle Prove: m<1 = m,2 + m<3 Shape will be provided below.

OpenStudy (anonymous):



OpenStudy (raden):

m1 and m4 is supplementary, so m1+m4=180degrees, or m1 = 180-m4 then the sum all angle of triangle is 180deg, gives m2+m3+m4=180 or m2+m3=180-m4 (remember just new, that 180-m4=m1) so, m2+m3=m1 proof

OpenStudy (anonymous):

Thank you!!

OpenStudy (raden):

welcome